# Deep Learning for Cross-Domain Data Fusion in Urban Computing: Taxonomy, Advances, and Outlook

### v1

@@ 3.2 Geographical Data
Geographical data plays a crucial role in modeling spatial relationships, contributing to the enhancement of cross-domain data fusion in urban computing by providing valuable insights into the geospatial context Breunig et al. (2020); Zhao et al. (2021a); Zhang et al. (2019). Tobler’s First Law of Geography, as stated by Miller (2004), indicates that ”Everything is related to everything else, but near things are more related than distant things.” This emphasizes the importance of geographical data, which serve as the basis for spatial modeling-based urban computing research. Figure 5 shows the main four types of geographical data through a visualization approach.
Points of Interest (POI) data, as a cornerstone of cross-domain data fusion in urban computing, command significant attention and utility in the realm of geographical data Lin et al. (2019); Wang et al. (2021a); Zhao et al. (2017); Balsebre et al. (2022); Fang et al. (2023); Cao et al. (2020); Gao et al. (2022c); Huang and Wang (2016). POI data refers to a collection of data representing specific locations or sites that hold significance or interest, often encompassing businesses, landmarks, or other notable entities in a given geographical area Psyllidis et al. (2022); Liu et al. (2020a). Conventionally, POI data may include the location coordinates (i.e., longitude and latitude), address, categorization (e.g., restaurants, hotels, parks, etc.), address, phone number, operating hours, and so on. Recent cross-domain urban datasets may also contain user reviews, photos, and other individual information Zhao et al. (2020); Chandra et al. (2021); Du et al. (2019); Balsebre et al. (2022). Therefore, POI data is capable of representing the semantics of a specific area, encapsulating key locations and entities to provide a comprehensive understanding of the geographic context. For example, Bing et al. (2023) collected POI data of two cities: New York City and Tokyo from the Foursquare platform and observed that POIs with higher spatial similarity values often have similar semantics. Through AMaps Service Platform in Wuhan and Shanghai, Bai et al. (2023) collected approximately 2 million POI records and fused them with satellite visual embedding for socioeconomic downstream tasks. To model the spatial correlations between POIs, Fu et al. (2019) developed a spatial network based on the multi-relation data among POIs, such as spatial distance and mobility connectivity.
@@
Satellite image data holds significant importance in representing geo-context by providing a visual depiction of Earth’s surface, and offering advantages such as global coverage, real-time monitoring, and the ability to capture fine-scale details Burke et al. (2021); Yuan et al. (2021b); Neupane et al. (2021). Satellite imagery is readily accessible and amenable to processing through various publicly available platforms (e.g., Google Earth Jenkins et al. (2019), OpenStreetMap Yin et al. (2023b), Baidu Map Yang et al. (2022); Xiao et al. (2023), and PlanetScope Li et al. (2022b)), facilitating multi-modal research in urban computing domain.
Moreover, street-view image data plays a vital role in depicting geo-contextual information by providing immersive, ground-level perspectives Biljecki and Ito (2021); Gong et al. (2019); Kang et al. (2020). A significant amount of cross-domain urban research has collected data from public platforms such as Google Street Liu et al. (2023g); Alfarrarjeh et al. (2021), Baidu Map Liu et al. (2023g); Jiang et al. (2021), and Tencent Map Huang et al. (2023b).
@@ 3.3 Traffic Data
Figure 6: Visualization of three primary types of traffic data: (a) traffic trajectory; (b) traffic flow; (c) road network Wu et al. (2016).
@@
Nowadays, location-based services, such as taxi and ride-hailing services, have generated enormous trajectory data. Additionally, many service providers, such as Uber and DiDi, have made their datasets publicly available to support research in this field. For instance, Zhang et al. (2021b) utilized taxi trip data during one month as a modality to assist in profiling urban regions. PANDA You et al. (2022) demonstrated the significance of taxi trajectory in predicting road risk, which can aid in the extraction of road segments. Geng et al. (2019b) forecasted ride-hailing demand on large-scale ride-hailing datasets in Beijing and Shanghai.
@@
Traffic flow data provides critical insights into the movement patterns of vehicles and pedestrians within urban environments, essential for understanding and optimizing city dynamics. In the realm of urban computing, three datasets — MobileBJ, BikeNYC, and TaxiBJ — emerge as key sources of insight for traffic mobility. Specifically, the MobileBJ dataset, collected from a Chinese social network, has significantly contributed to three key areas of urban computing. Firstly, it enabled the development of DeepSTN+ Lin et al. (2019), a deep learning model for predicting urban crowd flows, which integrates spatial dependencies, POIs, and temporal data. Additionally, it supported studies on high-dimensional spatio-temporal data in urban communities, employing advanced representation learning techniques Keerthi Chandra et al. (2020); Chandra et al. (2021).
@@
The BikeNYC dataset, derived from New York City’s bike-sharing system in 2014, provides rich data on trip durations, station locations, and times. It focuses on the last 14 days for testing purposes and includes 9 POI types. This dataset has been instrumental in developing a model for multi-step passenger demand forecasting Bai et al. (2019a), leveraging its detailed trip data to address complex spatio-temporal challenges in urban transportation planning Ji et al. (2023b); Qu et al. (2022); Zhang et al. (2023d). Meanwhile, TaxiBJ collects GPS traces from Beijing taxis, encompassing trip details, travel times, speeds, and specific pick-up and drop-off points. This dataset allowed for the development of a novel embedding strategy for urban regions Fu et al. (2019), providing deeper insights into city dynamics and supporting sustainable urban development. It also enhanced traffic prediction accuracy by applying spatiotemporal graph neural networks and integrating self-supervised learning for improved long-term forecasting Ji et al. (2023b). Furthermore, TaxiBJ demonstrated the effectiveness of a contrastive self-supervision method in inferring fine-grained urban flows, especially in environments with limited resources Qu et al. (2022). An innovative approach for computing trajectory similarities, TrajGAT, was introduced using TaxiBJ, significantly advancing the analysis of long trajectory data Yao et al. (2022).
@@
Road network data is intimately related to human daily lives, as it serves as the foundation for various services such as navigation and food delivery. The acquisition of road network data can be achieved through various methods. The earliest on-site manual surveying was labor-intensive and required a large amount of resources. With the development of remote sensing technology, which can effectively reduce costs, the proportion of road network data drawing relying on on-site collection has been decreasing. Another method to collect road network data is through UGC (User Generated Content), which collects road network information through anonymous terminal devices. The collected route can effectively help with updating road attributes and refinement of the road network. Generally, road network data can be downloaded in shapefile format from different open-sourced platforms such as Open Street Map, GRIP global roads database, DIVA-GIS, etc. For example, Yuan et al. (2021a) predicted travel demand and traffic flow based on taxi orders which is associated with trajectories on the road network. Zhu et al. (2020) ranked region significance taking into account multi-source spatial data including trajectories on road networks. Besides, the acquisition of street-view image datasets Huang et al. (2023a); Fan et al. (2023) also necessitated the sampling of collection points on road networks.
@@
Other miscellaneous data include logistic data, transportation safety data, and transportation recommendation data. In the logistics field, LaDe Wu et al. (2023b) introduced the first industry-scale last-mile delivery dataset. LaDe includes detailed information about the courier trajectories and waybill information, which can support massive spatio-temporal data mining tasks Ruan et al. (2022). Transportation safety data is also of great significance. For example, in Chen et al. (2018), road obstacle data was used to develop RADAR, a real-time system for identifying road obstacles in urban areas during typhoon seasons. You et al. (2022) used a comprehensive event dataset, including data on road accidents, fallen trees, and ponding water, to successfully develop a framework for predicting road risks in post-disaster urban settings with high accuracy. Transportation recommendations have become an integral part of our daily lives, contributing significantly to various service enhancements. For instance, Gao et al. (2023b) utilized this data to develop GraphTrip, a groundbreaking framework that leverages spatio-temporal graph representation learning for trip recommendations. This framework was rigorously tested using datasets from Edinburgh, Glasgow, Osaka, Toronto, and Melbourne, showcasing notable improvements in travel planning accuracy. Similarly, Guo et al. (2022) and He et al. (2019) employed multi-source urban data, encompassing operational data, taxi GPS trajectories, and public transportation information, to create advanced recommendation models.
@@ 3.4 Social Media Data
Twitter, an online social media and social networking service allows registered users to post geo-textual data Conger (2023). Hence, the Tweet data, owing to its inherent geo-tagged information, is utilized by researchers as a modality representing user social states, and it is integrated with other models in multi-modal learning. For instance, Zhao et al. (2016a) collected English tweets using the Twitter API in two cities, New York City and Singapore, and then annotated whether a tweet was POI-related. Finally, they can model the association between the tweet and its most semantically related POI. To study periodic human mobility patterns, Yuan et al. (2017) collected geo-annotated tweets from the most recent 3,200 tweets of Twitter users, and mapped them to the corresponding city by reverse geocoding. With a similar goal of understanding urban dynamics, Miyazawa et al. (2019) focused only on tweets regarding mobility and social activity; while Wang et al. (2015) are more interested in tweets concerning traffic events in Chicago. Furthermore, geo-textual data such as tweets can be used to learn user preferences to support personalized maps Zhao et al. (2016b). Some work, such as Vu et al. (2016) working on POI boundary estimation, also removed the content automatically created by other services like Twimight, Tweetbot, and so forth.
@@
Geo-tagged photos are crucial in multi-modal learning as they provide spatial context, enriching the understanding of content by incorporating location information into the broader spectrum of data modalities Nam and Yang (2022); Anbalagan (2022); Bui (2023). The two most commonly used geo-tagged photo datasets are the Yahoo Flickr Creative Commons (YFCC) dataset Thomee et al. (2015) as well as the NUS-WIDE dataset Chua et al. (2009). It is noteworthy that the YFCC dataset is well-known as the largest public multimedia collection released, which consists of 100 million photos posted on Flickr with relevant meta information such as geo-location coordinates and the date taken. The NUS-WIDE dataset is a well-known web image dataset (with 269,648 images and the associated tags from Flickr) created by Lab for Media Search in the National University of Singapore. Both Zhao et al. (2017) and He et al. (2019) leveraged the geo-tagged photos to jointly learn a context-aware embedding for personalized tour recommendation. Yin et al. (2019) derived training labels from the geo-tagged documents from the NUS-WIDE dataset, and therefore generated GPS embeddings. In the work of Yin et al. (2021) where multi context-aware location representations need to be learned, the geo-tagged photos from the NUS-WIDE dataset are used for image classification evaluation, whereas those from the YFCC dataset are leveraged to learn the semantic context.
@@
With the ubiquity of sensor-rich smartphones, acquiring continuous video frames with spatial metadata has become practical. The public geo-tagged mobile video data comes mainly from two mobile platforms, MediaQ Kim et al. (2014); MediaQ Project and GeoVid Arslan Ay et al. (2009); GeoVid Project . Lu et al. (2016) collected the geo-tagged video data from the aforementioned data platforms, and proposed the GeoUGV dataset consisting of two sets, videos and their geospatial metadata. Moreover, many social media platforms (e.g., WeChat Wang et al. (2021c), Gowalla Chen et al. (2015); Yuan et al. (2017), Baidu Liu et al. (2023b), Jiepang Fu et al. (2019)) encompass valuable user information, serving as a fundamental modality for fusion.
@@ 3.5 Demographic Data
The inclusion of demographic datasets in spatio-temporal multi-modal learning is pivotal as it enhances the contextual understanding of a given human group Song et al. (2019a); Wallin et al. (2019); Stratton (2021). The WorldPop organization plays a crucial role in global demographic research by providing high-resolution population data, enabling informed decision-making, and addressing various socio-economic and public health challenges worldwide Aheto et al. (2023); Luo et al. (2023); Cheng et al. (2023); Rogers et al. (2023). Xi et al. (2022) and Li et al. (2022b) collected Beijing population and population density statistics from the WorldPop platform as one of the predicted socioeconomic indicators; whereas Li et al. (2023d) extracted those from Singapore and New York City. In terms of data post-processing, Bai et al. (2023) further estimated the population density per grid cell through the Random Forest-based redistribution method.
Crime data holds immense significance as it serves as a critical resource for understanding patterns and factors influencing criminal activities, enabling policymakers and researchers to develop effective strategies for crime prevention and public safety Huang et al. (2018); Hajela et al. (2021); Zhang et al. (2022e). For example, Zhang et al. (2021b) collected crime data from the NYC Open Data website as ground truth values to be predicted by region embeddings, and there are 40 thousand crime records during one year in New York City.
Besides, land use data is crucial in urban planning, providing valuable insights into the spatial distribution of human activities, and informing decision-makers to optimize land resources for various purposes Steurer and Bayr (2020); Chen et al. (2021a); Risal et al. (2020). It consists of property data (e.g., private residential property transactions), residential data (e.g., buying and renting properties), business data (e.g., renewal of business use), and so on. For instance, Li et al. (2023d) collected land use data of Singapore and New York City from Singapore Master Plan 2019 and NYC MapPLUTO, respectively, as a region representation evaluation benchmark.
@@ 3.6 Environment Data
Environment data, particularly meteorological data, offers essential insights into dynamic weather patterns and environmental conditions that are integral for understanding complex interactions in various urban domains He et al. (2020); Anderegg et al. (2020); Perera et al. (2020); Trisos et al. (2020). The publicly available APIs can be utilized to gather meteorological data. For example, Yuan et al. (2021a) utilized the Dark Sky API integrated into Apple Weather to extract diverse weather characteristics for each region and used one-hot encoding to represent categorical attributes. You et al. (2022) extracted temporal contextual features including rainfall, temperature, humidity, dew point, and wind speed using Weather Underground API. Wang et al. (2021a) obtained temperature and sky condition data for both NYC and Chicago from NYC Open Data and Chicago Data Portal, respectively. By doing so, they were able to analyze the spatio-temporal correlations to predict the risk of traffic accidents. To precisely forecast the weather, Ma et al. (2023a) employed three real-world weather datasets: the WD_BJ dataset, which was gathered from 10 ground automatic weather stations in Beijing and contains nine meteorological variables; and the WD_ISR dataset and WD_USA dataset, both collected from OpenWeather and providing information on four weather conditions (namely temperature, humidity, wind speed, and atmospheric pressure) for Israel and the USA, respectively. In addition, enterprises also gather their own meteorological data, such as the DidiSY dataset which includes information on weather conditions, temperature, and wind speed specifically in Shenyang, a major city in China Bai et al. (2019b). Besides, open-source datasets from competitions can also be utilized, such as the Chinese weather dataset from the 7th Teddy Cup Data Mining Challenge, which includes 12 different types of weather Liu et al. (2022b).

### v2

@@ 3.2 Geographical Data
Geographical data plays a crucial role in modeling spatial relationships, contributing to the enhancement of cross-domain data fusion in urban computing by providing valuable insights into the geospatial context [21, 379, 377]. Tobler’s First Law of Geography, as stated by Miller [198], indicates that ”Everything is related to everything else, but near things are more related than distant things.” This emphasizes the importance of geographical data, which serve as the basis for spatial modeling-based urban computing research. Figure 6 shows the main four types of geographical data through a visualization approach.
Points of Interest (POI) data, as a cornerstone of cross-domain data fusion in urban computing, command significant attention and utility in the realm of geographical data [170, 272, 386, 13, 69, 31, 79, 107]. POI data refers to a collection of data representing specific locations or sites that hold significance or interest, often encompassing businesses, landmarks, or other notable entities in a given geographical area [224, 177]. Conventionally, POI data may include the location coordinates (i.e., longitude and latitude), address, categorization (e.g., restaurants, hotels, parks, etc.), address, phone number, operating hours, and so on. Recent cross-domain urban datasets may also contain user reviews, photos, and other individual information [380, 33, 63, 13]. Therefore, POI data is capable of representing the semantics of a specific area, encapsulating key locations and entities to provide a comprehensive understanding of the geographic context. For example, Bing et al. [20] collected POI data of two cities: New York City and Tokyo from the Foursquare platform and observed that POIs with higher spatial similarity values often have similar semantics. Through AMaps Service Platform in Wuhan and Shanghai, Bai et al. [10] collected approximately 2 million POI records and fused them with satellite visual embedding for socioeconomic downstream tasks. To model the spatial correlations between POIs, Fu et al. [74] developed a spatial network based on the multi-relation data among POIs, such as spatial distance and mobility connectivity.
@@
Satellite image data holds significant importance in representing geo-context by providing a visual depiction of Earth’s surface, and offering advantages such as global coverage, real-time monitoring, and the ability to capture fine-scale details [26, 353, 207]. Satellite imagery is readily accessible and amenable to processing through various publicly available platforms (e.g., Google Earth [116], OpenStreetMap [337], Baidu Map [324, 313], and PlanetScope [154]), facilitating multi-modal research in urban computing domain.
Moreover, street-view image data plays a vital role in depicting geo-contextual information by providing immersive, ground-level perspectives [18, 88, 130]. A significant amount of cross-domain urban research has collected data from public platforms such as Google Street [186, 4], Baidu Map [186, 124], and Tencent Map [113].
@@ 3.3 Traffic Data
Figure 7: Visualization of three primary types of traffic data: (a) traffic trajectory; (b) traffic flow; (c) road network [307].
@@
Nowadays, location-based services, such as taxi and ride-hailing services, have generated enormous trajectory data. Additionally, many service providers, such as Uber and DiDi, have made their datasets publicly available to support research in this field. For instance, Zhang et al. [368] utilized taxi trip data during one month as a modality to assist in profiling urban regions. PANDA [342] demonstrated the significance of taxi trajectory in predicting road risk, which can aid in the extraction of road segments. Geng et al. [84] forecasted ride-hailing demand on large-scale ride-hailing datasets in Beijing and Shanghai.
@@
Traffic flow data provides critical insights into the movement patterns of vehicles and pedestrians within urban environments, essential for understanding and optimizing city dynamics. In the realm of urban computing, three datasets — MobileBJ, BikeNYC, and TaxiBJ — emerge as key sources of insight for traffic mobility. Specifically, the MobileBJ dataset, collected from a Chinese social network, has significantly contributed to three key areas of urban computing. Firstly, it enabled the development of DeepSTN+ [170], a deep learning model for predicting urban crowd flows, which integrates spatial dependencies, POIs, and temporal data. Additionally, it supported studies on high-dimensional spatio-temporal data in urban communities, employing advanced representation learning techniques [134, 33].
@@
The BikeNYC dataset, derived from New York City’s bike-sharing system in 2014, provides rich data on trip durations, station locations, and times. It focuses on the last 14 days for testing purposes and includes 9 POI types. This dataset has been instrumental in developing a model for multi-step passenger demand forecasting [11], leveraging its detailed trip data to address complex spatio-temporal challenges in urban transportation planning [120, 226, 367]. Meanwhile, TaxiBJ collects GPS traces from Beijing taxis, encompassing trip details, travel times, speeds, and specific pick-up and drop-off points. This dataset allowed for the development of a novel embedding strategy for urban regions [74], providing deeper insights into city dynamics and supporting sustainable urban development. It also enhanced traffic prediction accuracy by applying spatiotemporal graph neural networks and integrating self-supervised learning for improved long-term forecasting [120]. Furthermore, TaxiBJ demonstrated the effectiveness of a contrastive self-supervision method in inferring fine-grained urban flows, especially in environments with limited resources [226]. An innovative approach for computing trajectory similarities, TrajGAT, was introduced using TaxiBJ, significantly advancing the analysis of long trajectory data [328].
@@
Road network data is intimately related to human daily lives, as it serves as the foundation for various services such as navigation and food delivery. The acquisition of road network data can be achieved through various methods. The earliest on-site manual surveying was labor-intensive and required a large amount of resources. With the development of remote sensing technology, which can effectively reduce costs, the proportion of road network data drawing relying on on-site collection has been decreasing. Another method to collect road network data is through UGC (User Generated Content), which collects road network information through anonymous terminal devices. The collected route can effectively help with updating road attributes and refinement of the road network. Generally, road network data can be downloaded in shapefile format from different open-sourced platforms such as Open Street Map, GRIP global roads database, DIVA-GIS, etc. For example, Yuan et al. [349] predicted travel demand and traffic flow based on taxi orders which is associated with trajectories on the road network. Zhu et al. [405] ranked region significance taking into account multi-source spatial data including trajectories on road networks. Besides, the acquisition of street-view image datasets [112, 68] also necessitated the sampling of collection points on road networks.
@@
Other miscellaneous data include logistic data, transportation safety data, and transportation recommendation data. In the logistics field, LaDe [305] introduced the first industry-scale last-mile delivery dataset. LaDe includes detailed information about the courier trajectories and waybill information, which can support massive spatio-temporal data mining tasks [235]. Transportation safety data is also of great significance. For example, in [39], road obstacle data was used to develop RADAR, a real-time system for identifying road obstacles in urban areas during typhoon seasons. You et al. [342] used a comprehensive event dataset, including data on road accidents, fallen trees, and ponding water, to successfully develop a framework for predicting road risks in post-disaster urban settings with high accuracy. Transportation recommendations have become an integral part of our daily lives, contributing significantly to various service enhancements. For instance, Gao et al. [78] utilized this data to develop GraphTrip, a groundbreaking framework that leverages spatio-temporal graph representation learning for trip recommendations. This framework was rigorously tested using datasets from Edinburgh, Glasgow, Osaka, Toronto, and Melbourne, showcasing notable improvements in travel planning accuracy. Similarly, Guo et al. [93] and He et al. [99] employed multi-source urban data, encompassing operational data, taxi GPS trajectories, and public transportation information, to create advanced recommendation models.
@@ 3.4 Social Media Data
Twitter, an online social media and social networking service allows registered users to post geo-textual data [50]. Hence, the Tweet data, owing to its inherent geo-tagged information, is utilized by researchers as a modality representing user social states, and it is integrated with other models in multi-modal learning. For instance, Zhao et al. [381] collected English tweets using the Twitter API in two cities, New York City and Singapore, and then annotated whether a tweet was POI-related. Finally, they can model the association between the tweet and its most semantically related POI. To study periodic human mobility patterns, Yuan et al. [352] collected geo-annotated tweets from the most recent 3,200 tweets of Twitter users, and mapped them to the corresponding city by reverse geocoding. With a similar goal of understanding urban dynamics, Miyazawa et al. [200] focused only on tweets regarding mobility and social activity; while Wang et al. [284] are more interested in tweets concerning traffic events in Chicago. Furthermore, geo-textual data such as tweets can be used to learn user preferences to support personalized maps [383]. Some work, such as [270] working on POI boundary estimation, also removed the content automatically created by other services like Twimight, Tweetbot, and so forth.
@@
Geo-tagged photos are crucial in multi-modal learning as they provide spatial context, enriching the understanding of content by incorporating location information into the broader spectrum of data modalities [203, 7, 25]. The two most commonly used geo-tagged photo datasets are the Yahoo Flickr Creative Commons (YFCC) dataset [263] as well as the NUS-WIDE dataset [47]. It is noteworthy that the YFCC dataset is well-known as the largest public multimedia collection released, which consists of 100 million photos posted on Flickr with relevant meta information such as geo-location coordinates and the date taken. The NUS-WIDE dataset is a well-known web image dataset (with 269,648 images and the associated tags from Flickr) created by Lab for Media Search in the National University of Singapore. Both Zhao et al. [386] and He et al. [99] leveraged the geo-tagged photos to jointly learn a context-aware embedding for personalized tour recommendation. Yin et al. [338] derived training labels from the geo-tagged documents from the NUS-WIDE dataset, and therefore generated GPS embeddings. In the work of Yin et al. [340] where multi context-aware location representations need to be learned, the geo-tagged photos from the NUS-WIDE dataset are used for image classification evaluation, whereas those from the YFCC dataset are leveraged to learn the semantic context.
@@
With the ubiquity of sensor-rich smartphones, acquiring continuous video frames with spatial metadata has become practical. The public geo-tagged mobile video data comes mainly from two mobile platforms, MediaQ [137, 197] and GeoVid [9, 85]. Lu et al. [187] collected the geo-tagged video data from the aforementioned data platforms, and proposed the GeoUGV dataset consisting of two sets, videos and their geospatial metadata. Moreover, many social media platforms (e.g., WeChat [277], Gowalla [42, 352], Baidu [172], Jiepang [74]) encompass valuable user information, serving as a fundamental modality for fusion.
@@ 3.5 Demographic Data
The inclusion of demographic datasets in spatio-temporal multi-modal learning is pivotal as it enhances the contextual understanding of a given human group [246, 271, 252]. The WorldPop organization plays a crucial role in global demographic research by providing high-resolution population data, enabling informed decision-making, and addressing various socio-economic and public health challenges worldwide [3, 189, 45, 233]. Xi et al. [309] and Li et al. [154] collected Beijing population and population density statistics from the WorldPop platform as one of the predicted socioeconomic indicators; whereas Li et al. [156] extracted those from Singapore and New York City. In terms of data post-processing, Bai et al. [10] further estimated the population density per grid cell through the Random Forest-based redistribution method.
Crime data holds immense significance as it serves as a critical resource for understanding patterns and factors influencing criminal activities, enabling policymakers and researchers to develop effective strategies for crime prevention and public safety [108, 95, 376]. For example, Zhang et al. [368] collected crime data from the NYC Open Data website as ground truth values to be predicted by region embeddings, and there are 40 thousand crime records during one year in New York City.
Besides, land use data is crucial in urban planning, providing valuable insights into the spatial distribution of human activities, and informing decision-makers to optimize land resources for various purposes [250, 35, 231]. It consists of property data (e.g., private residential property transactions), residential data (e.g., buying and renting properties), business data (e.g., renewal of business use), and so on. For instance, Li et al. [156] collected land use data of Singapore and New York City from Singapore Master Plan 2019 and NYC MapPLUTO, respectively, as a region representation evaluation benchmark.
@@ 3.6 Environment Data
Environment data, particularly meteorological data, offers essential insights into dynamic weather patterns and environmental conditions that are integral for understanding complex interactions in various urban domains [101, 8, 222, 266]. The publicly available APIs can be utilized to gather meteorological data. For example, Yuan et al. [349] utilized the Dark Sky API integrated into Apple Weather to extract diverse weather characteristics for each region and used one-hot encoding to represent categorical attributes. You et al. [342] extracted temporal contextual features including rainfall, temperature, humidity, dew point, and wind speed using Weather Underground API. Wang et al. [272] obtained temperature and sky condition data for both NYC and Chicago from NYC Open Data and Chicago Data Portal, respectively. By doing so, they were able to analyze the spatio-temporal correlations to predict the risk of traffic accidents. To precisely forecast the weather, Ma et al. [192] employed three real-world weather datasets: the WD_BJ dataset, which was gathered from 10 ground automatic weather stations in Beijing and contains nine meteorological variables; and the WD_ISR dataset and WD_USA dataset, both collected from OpenWeather and providing information on four weather conditions (namely temperature, humidity, wind speed, and atmospheric pressure) for Israel and the USA, respectively. In addition, enterprises also gather their own meteorological data, such as the DidiSY dataset which includes information on weather conditions, temperature, and wind speed specifically in Shenyang, a major city in China [12]. Besides, open-source datasets from competitions can also be utilized, such as the Chinese weather dataset from the 7th Teddy Cup Data Mining Challenge, which includes 12 different types of weather [178].